A carpenter charges a set fee for each job in addition to his hourly charge. If a carpenter bills a customer $150 for a job which takes him 3 hours and bills a second customer $180 for a 6 hour job, how much would he charge for a job that takes him 9 hours?
Answer by askmemath(368)

Let his hourly cost = c
Then for 3 hours he will charge 3c
So his total cost = X+3c where X is his fixed fee
X+3c = 150
X+ 6c = 180
Subtracting the first eqn from the second we get 3c = 30 or c= 10
Substituting in any equation we get X = 120
So for a 9 hour job he will charge X + 9c = 120 +9x10 = 120+90 = $210